OpenStudy (anonymous):

Is the answer $74.25? The original value of a car is $22,000, and it depreciates (loses value) by 15% each year. What is the value of the car after three years? $15,895 $13,510.75 $56,595 $74.25 22000, 3300, 495, 74.25 Does it mean the third term or the forth? That confused me. 495 isn't an answer choice so I kept going...

OpenStudy (ranga):

The value decreases by 15% each year. This is same as saying the value will be 0.85 times the original value after each year. After first year the value will be: 22,000 * .85 = 18,700 After second year the value will be: 18,700 * .85 = 15,895 After third year the value will be: 15,895 * .85 = $13,510.75

OpenStudy (anonymous):

wait were did you get .85 from?

OpenStudy (ranga):

The value decreases by 15% each year. So if I start with, say, $100, the value will drop by 100 * 15 / 100 = $15 after one year. So what was $100 now will be 100 - 15 = $85 after one year which is 0.85 * $100. So 15% reduction in value each year is same as multiplying by 0.85 each year

OpenStudy (ranga):

If you want to do it the longer route you can. Start with $22,000 A 15% reduction is: 22000 * 15 / 100 = $3,300 So the value after one year will be $22,000 - $3,300 = $18,700 We could have quickly arrive at the same result by multiplying 22000 by 0.85 = 18,700

OpenStudy (ranga):

And you have to continue the above for two more years to get the final answer of $13,510.75

OpenStudy (anonymous):

thank you :)

OpenStudy (ranga):

You are welcome.